Lehighton Chalk Company manufactures sidewalk chalk, which it sells online by the box at $22 per unit. Lehighton uses an actual costing system, which means that the actual costs of direct material, direct labor, and manufacturing overhead are entered into work-in-process inventory. The actual application rate for manufacturing overhead is computed each year; actual manufacturing overhead is divided by actual production (in units) to compute the application rate. Information for Lehighton’s first two years of operation is as follows:

Year 1 Year 2
Sales (in units) 3,100 3,100
Production (in units) 3,600 2,600
Production costs:
Variable manufacturing costs $ 15,840 $ 11,440
Fixed manufacturing overhead 19,440 19,440
Selling and administrative costs:
Variable 12,400 12,400
Fixed 11,400 11,400

Selected information from Lehighton’s year-end balance sheets for its first two years of operation is as follows:

LEHIGHTON CHALK COMPANY
Selected Balance Sheet Information
Based on absorption costing End of Year 1 End of Year 2
Finished-goods inventory $ 4,900 $ 0
Retained earnings 8,520 14,440
Based on variable costing End of Year 1 End of Year 2
Finished-goods inventory $ 2,200 $ 0
Retained earnings 5,820 14,440
  1. Reconcile Lehighton’s operating income reported under absorption and variable costing, during each year, by comparing the following two amounts on each income statement:

  • Cost of goods sold
  • Fixed cost (expensed as a period expense)
  1. What was Lehighton’s total operating income across both years under absorption costing and under variable costing?

  2. What was the total sales revenue across both years under absorption costing and under variable costing?

  3. What was the total of all costs expensed on the operating income statements across both years under absorption costing and under variable costing?

  4. Subtract the total costs expensed across both years [requirement (4)] from the total sales revenue across both years [requirement (3)]: (a) under absorption costing and (b) under variable costing.

  5. Considering the results obtained in requirements 1-5 above, select which of the following statements (is) are true by selecting an "X".
Answer #1
ncome Statement - Absorption Cosing
Particulars Year 1 Year 2
Details Amount Details Amount
Unit 3100 3100
Sales $68,200.00 $68,200.00
Cost of Goods Sold:
Cost of goods produced $35,280.00 $30,880.00
Add: Opening Inventory $0.00 $4,900.00
Less: Ending Inventory $4,900.00 $30,380.00 $0.00 $35,780.00
Gross Profit $37,820.00 $32,420.00
Variable Selling & Administrative Expenses $12,400.00 $12,400.00
Fixed Selling & Administrative Expenses $11,400.00 $11,400.00
Net Operating Income $14,020.00 $8,620.00
Variable costing contribution format income statement
Particulars Year 1 Year 2
Details Amount Details Amount
Sales 3100*$22 $68,200.00 3100*$22 $68,200.00
Variable Cost:
Variable manufacturing cost (15840/3600*3100) $13,640.00 (11440/2600*3100) $13,640.00
Variable Selling and Administrative Expenses $12,400.00 $12,400.00
Contribution $42,160.00 $42,160.00
Fixed Manufacturing Overhead $19,440.00 $19,440.00
Fixed Selling & Administrative Expenses $11,400.00 $11,400.00
Net Income $11,320.00 $11,320.00
Sales Revenue as per Absorption Costing & Variable costing
Absroption Costing Variable costig
Year-1 (3100*22)=$68200 (3100*22)=$68200
Year-2 (3100*22)=$68200 (3100*22)=$68200
Cost expensed as per Absorption Costing & Variable costing
Absroption Costing Variable costig
Year-1 (30380+23800)=$54180 (26040+30840)=$56880
Year-2 (35780+23800)=$59580 (26040+30840)=$56880
Part-4
Absroption Costing Variable costig
Year-1 Year-2 Year-1 Year-2
Revenue $68,200.00 $68,200.00 $68,200.00 $68,200.00
Cost $54,180.00 $59,580.00 $56,880.00 $56,880.00
Operating Income $14,020.00 $8,620.00 $11,320.00 $11,320.00